OpenStudy (austinlr12):

AB and BC form a right angle at point B. If A = (-3, -1) and B = (4, 4), what is the equation of BC? x + 3y = 16 2x + y = 12 -7x − 5y = -48 7x − 5y = 48

OpenStudy (austinlr12):

@Michele_Laino

OpenStudy (austinlr12):

@AdoNine

OpenStudy (michele_laino):

here, we have to write the equation of line AB first

OpenStudy (adonine):

yes

OpenStudy (michele_laino):

the slope \(m\) of such line is: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 1 - 4}}{{ - 3 - 4}} = ...?\]

OpenStudy (austinlr12):

5/7?

OpenStudy (austinlr12):

Would the answer be 7x − 5y = 48 ?

OpenStudy (michele_laino):

correct! I think it is suffice so since line BC is perpendicular to line AB, then its slope \(m'\) has to check this identity: \(m \cdot m'=-1\) namely: \[m' = \frac{{ - 1}}{m} = \frac{{ - 1}}{{5/7}} = ...?\]

OpenStudy (adonine):

@austinlr12 you are correct

OpenStudy (austinlr12):

Thxs

OpenStudy (michele_laino):

Please wait, I'm checking... and the requested equation, is therefore: \[y - 5 = \frac{{ - 7}}{5}\left( {x - 5} \right)\]

OpenStudy (michele_laino):

oops.. I have made a typo: \[y - 4 = \frac{{ - 7}}{5}\left( {x - 4} \right)\]

OpenStudy (michele_laino):

I got this equation: \[7x + 5y = 48\]

OpenStudy (austinlr12):

I already submit the answer...